OpenStudy (anonymous):

Without dividing, do you expect five-eighths divided by one-sixteenth to be greater than or less than 1? Explain.

OpenStudy (anonymous):

Remember if you're dividing a fraction by a fraction, the denominator (the number of bottom) of the bottom fraction gets flipped up to the very top \[ \frac{\frac{5}{8}}{\frac{1}{16}} = \frac{5}{8} \frac{16}{1} \] And then if the numerator is bigger than the denominator the fraction is greater than 1 :) Follow?

OpenStudy (anonymous):

that means that it would be greater than 1 right?

OpenStudy (anonymous):

yah! :D

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

my pleasure